OpenStudy (anonymous):

Two lines, A and B, are represented by the following equations: Line A: 3x + 3y = 12 Line B: x + y = 4 Which statement is true about the solution to the set of equations? It is (12, 4). There are infinitely many solutions. It is (4, 12). There is no solution. plz plz plz help me

OpenStudy (butterflydreamer):

Note: you can factorise the equation for line A. 3x + 3y = 12 factorise the 3 and we get: 3 ( x + y) = 12 Then look at the equation for line B : x + y = 4 So we can know "x + y = 4"... Plug this into line A " 3 (x + y ) = 12" 3(4) = 12 12 = 12 Therefore, this means that there will be infinitely many solutions.

OpenStudy (anonymous):

ty sm :)